Constants Part A At the instant the traffic light turns green, a car that has been waiting at an intersection starts ahead with a constant acceleration of 2.80 m/s2. At the same instant a truck, traveling with a constant speed of 20.0 m/s, overtakes and passes the car How far beyond its starting point does the car overtake the truck? Express your answer with the appropriate units. FalueUnits Submit ▼ Part B How fast is the car traveling when it overtakes the truck? Express your answer with the appropriate units IA vValue Units Submit Part C Complete previous part(s) Part D Complete previous part(s)
media%2Fca8%2Fca830989-b35d-4bef-b6b6-c0
media%2F980%2F980f6154-772d-49d0-8aaf-5e
Answer #1

A)

vo = initial velocity of the car = 0 m/s

a = acceleration of the car = 2.80 m/s2

V = velocity of truck = 20 m/s

let the car catch the truck after time "t"

x = displacement of the car in time "t"

displacement of the car in time "t" is given as

x = vo t + (0.5) a t2                                             eq-1

X = displacement of truck in time "t"

displacement of the truck in time "t" is given as

X = V t                                                                    eq-2

for the car to catch the truck ,

x = X

using eq-1 and eq-2

vo t + (0.5) a t2 = V t

(0) t + (0.5) (2.80) t2 = 20 t

t = 14.3 sec

using eq-2

X = V t   

X = 20 x 14.3 = 286 m

b)

velocity of the car is given as

v = vo + a t

v = 0 + (2.80) (14.3)

v = 40 m/s